OpenStudy (anonymous):

please help im struggling please explain to me the answer 1. Part A.) For the f(x) = 5x + 8 and g(x) = x – 3. Calculate (f o g)(x) Part B.) f(x) = 3x + 5 and g(x) = x². Calculate (f o g)(-3)

OpenStudy (tkhunny):

This is the principle... f(x) = ax+b f(q) = a(q) + b f(Fred) = a(Fred) + b f(Pittsburgh) = a(Pittsburgh) + b So, what do you suppose f(g(x)) would be?

OpenStudy (anonymous):

i have no idea im just really confused

OpenStudy (tkhunny):

That's not really good enough. I just gave you three examples. Please apply what you see there to answer the question. You may NOT have "no idea".